OpenStudy (anonymous):

how do i solve the equation : 5/12=10/x-1

OpenStudy (anonymous):

x=25

OpenStudy (anonymous):

5/12=10/x-1 by cross multiplication 5(x-1)=10(12) 5x-5=120 5x=125 x=25